OpenStudy (liizzyliizz):

find the solution to dT-k(T+45)dt=0 that satisfies T(0)=200

OpenStudy (anonymous):

the way you are writing things out is confusing !!!!

OpenStudy (liizzyliizz):

That is exactly how it is written... and I have it.

OpenStudy (liizzyliizz):

I am lost too..

OpenStudy (amistre64):

i believe its separable

OpenStudy (liizzyliizz):

yeah my first step would be to separate variables and then integrate.

OpenStudy (liizzyliizz):

I am assuming this is just another differential equation that I am solving,

OpenStudy (amistre64):

\[dT-k(T+45)dt=0\] \[dT=k(T+45)dt\] \[(T+45)^{-1} dT=k\ dt\]

OpenStudy (amistre64):

yes it is

OpenStudy (amistre64):

we might even be able to do it other ways; but thats the first thing that came to mind

OpenStudy (liizzyliizz):

quick question where did that -1 come from o.o

OpenStudy (amistre64):

\[\frac{1}{a}=a^{-1}\] i just divided both sides by T+45

OpenStudy (amistre64):

\[\frac{1}{T+45}dT=k\ dt\] same thing

OpenStudy (liizzyliizz):

ahh ok :) got cha,